A gas at a volume in a container at pressure is compressed to on-fifth of it's volume. What will it's pressure be if it's temperature remains constant


1
Expert's answer
2021-05-17T04:20:31-0400

P1 = Po

V1 = Vo

T1 = T2 = const

P2 = unknown

V2 = (1/5)Vo = 0.2Vo


Solution:

Since the temperature and amount of gas remain unchanged, Boyle's law can be used.

Boyle's gas law can be expressed as: P1V1 = P2V2

To find the new pressure, solve the equation for P2:

P2 = P1V/ V2

P2 = (Po × Vo) / (0.2Vo) = 5Po

P2 = 5Po

Thus, if the volume is decreased 5 times, the pressure will increase 5 times.


Answer: 5Po